OpenStudy (anonymous):

PLEASE HELP. Write the sum using summation notation, assuming the suggested pattern continues. -1 + 2 + 5 + 8 + ... + 44

OpenStudy (amistre64):

so whats our pattern?

OpenStudy (anonymous):

+3 I'm mainly confused on how to set up the summation notation.

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

well if we have a rule, then we can just tell the summation to add up all the terms defined by the rule we start at -1: and add 3 soo -1 +3n is a good rule, for n=0,1,2,3,...

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

if we want to sum up all the terms, \[\Sigma\text{; is just a capital S, for 'summation"}\] \[\sum_{n=low}^{high}rule\]

OpenStudy (amistre64):

instead of greek :) \[\{Sum\}_{n=0}^{n=k}~~~(rule)\]

OpenStudy (anonymous):

sooo \[\sum_{n=0}^{15} (-1+3n) ?????\] or would it be infinity instead of 15..??

OpenStudy (amistre64):

hmm, i spose we should try to figure out how many terms we have ...

OpenStudy (amistre64):

-1 + 3n = 44 what is n?

OpenStudy (amistre64):

45/3 = 15 yeah, thats fine :)

OpenStudy (anonymous):

okay ^.^

OpenStudy (anonymous):

so I'm right??

OpenStudy (amistre64):

first term: n=0, we get -1 last term: n=15, we get 44 thats all of them

OpenStudy (amistre64):

yep, you did great

OpenStudy (anonymous):

yay! :D

OpenStudy (anonymous):

can you help me with more?? i suck at this stuff... >.<

OpenStudy (amistre64):

one more, maybe

OpenStudy (anonymous):

thanks i'll make it a good one lol

OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. 4 - 24 + 144 - 864 + ...

OpenStudy (amistre64):

alternating signs eh, whats that suggest?

OpenStudy (anonymous):

multiplication is my best guess....

OpenStudy (amistre64):

yep, and (-1)^n so that the signs alternate now what our ratio?

OpenStudy (anonymous):

-6...? maybe.. i'm not confident in this answer...

OpenStudy (amistre64):

your right, dont worry (-6)^n = (-1)^n 6^n so it all depends on how you want to present that part of it

OpenStudy (amistre64):

then n=0, (-6)^n = 1 and we multiply our first term so whats our rule?

OpenStudy (amistre64):

consider 4 = a (-6)^0 what is a?

OpenStudy (anonymous):

the rule will always be the first term plus the ratio right??

OpenStudy (amistre64):

well, times the ratio in this case. its common differences that add, ratios multiply

OpenStudy (amistre64):

a0 = a a1 = ar a2 = arr a3 = arrr each new term is the last one times r

OpenStudy (anonymous):

that confused me more...

OpenStudy (anonymous):

4(-6)^n-1??

OpenStudy (amistre64):

does n start at 1 or 0?

OpenStudy (anonymous):

i'm not sure.... but im guessing 0 so it would be... 4(-6)^n

OpenStudy (amistre64):

its author dependant, i start at n=0, some courses say n=1 so there is really no standard to go by. as long as we define n starts at 0 it makes the rule simpler to write

OpenStudy (amistre64):

now we hav eno final term, so what our upper limit?

OpenStudy (amistre64):

so, write up the summation notation :)

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty} 4(-6)^n \] ...?

OpenStudy (amistre64):

perfect

OpenStudy (anonymous):

YAY :D Thank you soooooooooo much!!!!!!

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

thanks! :)